OpenStudy (anonymous):

HELP ME PLEASE!!! Divide3K^2/4K-2 / K^2+3K+2/2K^2+5K-3

OpenStudy (anonymous):

@timo86m help?

OpenStudy (anonymous):

is your question divide \[\frac{ 3k ^{2} }{4k-2 } by \frac{ k ^{2}+3k+2 }{2k ^{2}+5k-3 } ~?\]

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

I think that is what she meant surji. Use perenthesis next time pinky. and first step is to take the reciprocal of the second ratio :D

OpenStudy (anonymous):

http://openstudy.com/users/timo86m#/updates/52fe7476e4b0332f69f7dec3 THis is similar pink mommy

OpenStudy (anonymous):

that whole thing confuses me.

OpenStudy (anonymous):

Can you show me how to do this one step by step please?

OpenStudy (anonymous):

well first of all use perenthesis and spaces makes it easier to read ((3K^2)/(4K-2)) / ((K^2+3K+2)/(2K^2+5K-3)) here is a template ( ( )/( ) ) / ( ( )/( ) )

OpenStudy (anonymous):

when in doubt use perenthesis for groupings

OpenStudy (anonymous):

@timo86m sorry I didn't know how to write it correctly.

OpenStudy (anonymous):

\[\frac{ 3k ^{2} }{ 2\left( 2k-1 \right)}\times \frac{ 2k ^{2}+5k-3 }{ k ^{2}+3k+2 }\] make factors and then cancel if you have common factors.

OpenStudy (anonymous):

It's ok but it is very important :) cuzz if you know how then you can use wolfram. It jsut makes it able for computers and humans to read it :)

OpenStudy (anonymous):

((3K^2)/(4K-2)) / ((K^2+3K+2)/(2K^2+5K-3)) first thing is to take reciprocal of second ratio: I do it below and notice the / is now a * :) ((3K^2)/(4K-2)) * ((2K^2+5K-3)/(K^2+3K+2)) i will try is to factor them into binomials leaving noe k^2 but only ks: as shown below ((3K^2)/(4K-2)) * ((2K^2+5K-3)/(K^2+3K+2)) ^ ^ already factored ^ ^ try it out :)

OpenStudy (anonymous):

if yo ucant factor it leave it alone and move on to next one. Like 3k^2 i think you cant so move on to the next one :)

OpenStudy (anonymous):

(k+3)(2k−1) /(K+2)(K+1)

OpenStudy (anonymous):

2*K^2+5*K-3 = (K + 3) (2 K - 1) K^2+3*K+2 = (K + 2) (K + 1) try to them them cancel :)

OpenStudy (anonymous):

nothing to cancel

OpenStudy (anonymous):